# Train Once, Deploy Anywhere: Realize Data-Efficient Dynamic Object Manipulation

### v1

@@ V-E4 Generalization across Robot Embodiments
Franka Franka 0.88 0.98 0.83 0.80
Franka UR10e 0.91 0.97 0.85 0.79
Piper Piper 0.92 0.96 0.79 0.74
Piper ARX5 0.90 0.95 0.77 0.69
TABLE XIII: Analysis on sim-to-real generalization.
@@
$\times$ 0.76 0.83 0.55 0.48
$\checkmark$ 0.92 0.96 0.79 0.74
V-E5 Sim-to-Real Generalization
The input to the recurrent probabilistic network is uncolored 3D points, which do not exhibit significant differences between simulated and real environments. Therefore, GEM solely based on simulated demonstrations can generalize to the real world to perform operations. To validate this characteristic, we report the real-world manipulation success rates of GEMs with and without tuning in Table XIII. According to the results, we can observe that the GEM trained using only simulated data can manipulate objects in the real world, though its success rates are lower than the GEM tuned with 20 demonstrations per skill. Through analysis, we find this is caused by two reasons. First of all, without tuning, the test objects are all unseen in demonstrations, which increases the manipulation challenge. Secondly, in a robot control system, there is a control execution delay between the input signal and the system response. This control execution delays are different between simulated and real robots, further enhancing the difficulty.
V-F Real Application Deployment
@@
Since onsite data collection is infeasible in this canteen, we teleoperate 100 demonstrations using the manipulation platform shown in Fig. 5. Then, we utilize these 100 demonstrations to fine-tune the GEM policy that has been pre-trained with the 6,600 simulated data. Finally, we deploy the fine-tuned GEM in this canteen to conduct spoon collection. We evaluate GEM over a seven-day period, processing more than ten thousand spoons, and measure a final manipulation success rate of 97.2%. This marks the first real application of imitation learning in dynamic object manipulation.
@@ VI Conclusion
This paper has first developed an entropy-based theoretical framework to analyze the data requirements of imitation learning. Based on this framework, we have proposed GEM that reduces observation complexity and action prediction ambiguity to achieve strong generalization. Through extensive experiments, GEM has successfully generalized across diverse environment backgrounds, motion patterns, object geometries, and robot embodiments, achieving a success rate of over 97% in a real-world application without using any in-scene demonstration. These results have demonstrated that our theoretical insights effectively guide the design of an efficient manipulation system, offering a practical solution for industrial automation with reduced data costs.

### v2

@@ V-D4 Generalization across Robot Embodiments
Franka Franka 0.94 1.00 0.90 0.86
Franka UR10e 0.95 1.00 0.88 0.86
Piper Piper 0.98 1.00 0.93 0.89
Piper ARX5 0.96 1.00 0.94 0.85
TABLE XIII: Sim-to-real performance of GEM.
@@
Real_tune 0.98 1.00 0.93 0.89 0.95
Real_zero 0.91 0.98 0.80 0.79 0.87
V-D5 Sim-to-Real Generalization
According to the aforementioned experiments, GEM can generalize to novel environment backgrounds, motion patterns, target objects, and robot embodiments. These characteristics enable the possibility of generalizing to real-world tasks based on simulated demonstrations. To validate this issue, we design two real-robot experiment settings, i.e., tuned sim-to-real and zero-shot sim-to-real. The experimental results of the tuned sim-to-real setting are presented in Table II. The results suggest that GEM achieves significantly superior sim-to-real generalization with a handful of tuning demonstrations.
The performance of GEM in the zero-shot sim-to-real setting is reported in Table XIII. According to the results, it is observed that the GEM policy trained using only simulated data can manipulate objects in the real world, and its success rates are still promising. This strong generalization is attributed to the design that GEM prioritizes geometry information and decreases the importance of visual appearance. We also analyze the reasons of the slight performance deterioration. An important factor is there exists a control execution delay between the input signal and the robot system response. This control execution delays are different between simulated and real robots, causing a sim-to-real execution gap. Tuning the policy with a few real-robot demonstrations can bridge this execution delay gap. Therefore, the performance of GEM in the tuned sim-to-real setting is stronger.
V-E Real Application Deployment
@@
Since onsite data collection is infeasible in this canteen, we teleoperate 100 demonstrations using the manipulation platform shown in Fig. 8. Then, we utilize these 100 demonstrations to fine-tune the GEM policy that has been pre-trained with the simulation demonstrations. Finally, we deploy the fine-tuned GEM in this canteen to conduct spoon collection. We evaluate GEM over a seven-day period, processing more than ten thousand spoons, and measure a final manipulation success rate of 97.2%.
@@ VI Conclusion
This paper has first pointed out that embodied action generation concerns more about the 3D geometry structure of environments than visual appearance. Based on this insight, we have proposed the policy GEM for dynamic object manipulation on conveyor systems. The key design principle of this policy is prioritizing the exploitation of 3D geometry information and decreasing the reliance on 2D visual information. In addition, we have developed simulation platforms of the explored tasks to generate training demonstrations efficiently. Since the simulators can simulate the geometry-based interaction dynamics among objects well and GEM primarily uses geometry information, GEM achieves promising sim-to-real generalization. We have conducted extensive experiments to analyze the characteristics of GEM. The results indicate that GEM has successfully generalized across diverse environment backgrounds, motion patterns, object geometries, and robot embodiments, achieving a success rate of over 97% in a real-world application without using any in-scene demonstration. These results have demonstrated that our policy design principle effectively guides the design of an efficient manipulation system, offering a practical solution for industrial automation.